OpenStudy (anonymous):

A television game has 6 shows doors, of which the contests must pick 2.behind two of the doors are expensive cars, and behind the other 4 doors are consolation prizes. Find the probl that the contestant wins exactly 1 car, no car, or atleast one car

OpenStudy (anonymous):

Find the prob the contest wins exacly one car

OpenStudy (anonymous):

there are 6 choose 2 ways to pick 2 doors. 6 choose 2 is 6*5/2 = 15. this is your denominator. there are 8 ways to pick one car and one consolation prize. answer to the first question is 8/15 there are 4 choose 2 ways to pick 2 out of the 4 consolation prizes. 4 choose 2 is 6. answer to the second question is 6/15. at least one car is the same as 'not zero cars' so that probability is 1 - 6/15 = 9/15 = 3/5